2 whole 2/3 ÷ 5 is equal to (a) 8/15 (b) 40/3 (c) 40/5 (d) 8/3

Grade:12

1 Answers

Harshit Singh
3 years ago
Dear Student

Option (a) is correct here

First we have to convert the mixed fraction into improper fraction 2 whole 2/3 = 8/3 Then,

= (8/3) ÷ 5
= (8/3)/(5/1)
= (8/3) × (1/5)
= 8/15


Thanks